# Stability of Ecological SystemsCan Chen, Xu-Wen Wang, and Yang-Yu Liu

### 12 Dec 2023 v1

@@ 3.2.2 Impacts of extinction
Extinction is generic in the GLV model with random interactions [38, 88, 195]. Recent studies have highlighted that an equilibrium point of the GLV model is locally stable if and only if all the eigenvalues of the reduced interaction matrix (i.e., the interaction matrix between the species in the surviving sub-community) have negative real parts [31, 38, 53, 244]. Baron et al. focused on the eigenvalue spectrum of the reduced interaction matrix in the GLV model, where the spectrum of the reduced interaction matrix also consists of a bulk set of eigenvalues and an outlier [38]. Importantly, the authors demonstrated that the universality principle holds for the bulk region (i.e., following Ellipse Law) but not for the outlier eigenvalue. The outlier eigenvalue can be solved from the generating functional approach. Prediction of extinction boundary and new equilibrium after a primary extinction were also discussed in [195] and [88], respectively.
Figure 5: Eigenvalue distributions with time delay, where the blue teardrop-shaped area represents the stability region with time delay. (a) The eigenvalues (represented by red dots) are within the teardrop-shaped area, indicating that the corresponding GLV model with time delay is locally stable. (b) The eigenvalues are outside the teardrop-shaped area, indicating that the corresponding GLV model with time delay is unstable. $\tau=1$. The figure was redrawn from [271].
@@ 3.2.3 Impacts of time delay
Similar to Pigani et al.’s work as discussed in Section 3.1.7, Yang et al. investigated the stability of the time-delayed GLV model [271], which is defined as
@@
where $\textbf{M}_{\text{delay}}=\textbf{X}\textbf{A}$ is the community matrix with delay. For simplicity, X is set to I such that each species has unit abundance. However, the validation of this assumption remains an open problem [271]. They proved that (12) is stable if all the roots of the characteristic equation, defined as $H(z)=z-\lambda_{\text{delay}}e^{-z\tau}$ where $\lambda_{\text{delay}}$ are the eigenvalues of $\textbf{M}_{\text{delay}}$, have negative real parts. This equation implies that all the eigenvalues of $\textbf{M}_{\text{delay}}$ are required to be located in a teardrop-shaped region defined by $\text{Re}[z]<0$ to ensure stability, see Fig. 5. Since the distribution of the eigenvalues of $\textbf{M}_{\text{delay}}$ (i.e., the interaction matrix A is this case) with random, mutualistic, or competitive interactions is well-established, it becomes straightforward to estimate $\max{(\text{Re}[z])}$. Based on the theoretical findings, the authors found that time delay plays a critical role in ecological systems, where large delay is often destabilizing, but short delay can considerably enhance community stability ($\tau=0.2$ and 0.5 were used for the small and large time delay in their simulations, respectively) [271].
Similar conclusions were also obtained by Saeedian et al. when investigating the impact of time delay on the emergent stability patterns within the time-delayed GLV model (11) [222]. Significantly, they further determined the existence of a Hopf bifurcation (i.e., the two complex conjugate eigenvalues of the community matrix, with non-zero imaginary part, simultaneously cross the imaginary axis into the right half-plane) at $\tau_{c}$, which is computed as
@@ 3.2.4 Impacts of stochastic noises
Real ecological systems are inherently stochastic with constant external perturbations and internal fluctuations [252]. Krumbeck et al. developed a theoretical framework for the analysis of temporal stability of ecological systems with stochastic noises [146]. They provided analytical predictions for the power spectral density of the stochastic GLV model defined as
@@
where $i$ here is the imaginary number. They showed that different network structures (random, mutualistic, and competitive) have unique signatures in the spectrum of fluctuations. They further investigated the effect of trophic structures and identified a gap in the power spectral density, which indicates that high-level trophic structures contribute to enhanced long-term temporal stability [146].
@@ 3.2.5 Impacts of evolved system size
Previous work on the linear stability of the GLV model is concerned with ecological systems of predetermined and fixed sizes. Galla explored the stability of ecological systems with evolved system size using methods from statistical mechanics and the theory of disordered systems [96, 161, 169]. Specifically, he exploited generating functionals to derive the effective dynamics for the GLV model (with $r_{i}=1$) computed as
@@
This result implies that predator-prey relationships enhance stability, while variability in species interactions promotes instability [96], which aligns with prior findings as reported in [249, 270].
Poley et al. applied Galla’s approach to analyze the stability of the GLV model with hierarchical interactions [203] defined as
@@
for $i=1,2,\dots,S^{a}$, where $B$ is the total number of local systems, and $a,b=1,2,\dots,B$ denote the indices of the local systems with the associated system size $S^{a}$ and $S^{b}$. Their findings indicate that a strong hierarchical structure is stabilizing but reduces diversity within the effective dynamics of the GLV model [203]. Similarly, Sidhom and Galla adapted Galla’s approach to study the GLV model with nonlinear feedback [229] defined as
@@
for $i=1,2,\dots,S$, where the function $g(u)=\frac{2au}{a+2|u|}$ denotes the nonlinear feedback ($a$ is the saturation parameter). This form of feedback was initially introduced to describe the predator’s growth rate when interacting with prey. It is plausible that the benefits from extra prey will eventually saturate as prey numbers become large. The authors concluded that the stability and diversity of ecological systems improves with the introduction of nonlinear feedback [229].
Figure 6: Stability of real and permuted food webs in relation to complexity. These permuted food webs were constructed from real food webs by removing some non-random features, namely the row structure, topology, pairwise correlation, and interaction strength. The more the “rightmost”eigenvalues, i.e., $\text{Re}[\lambda_{1}]$, close to zero, the more stable of the food webs (because the diagonal elements of the community matrix are set to zero). This figure was redrawn from [126].
@@ 3.2.6 Applications to real ecological systems
Last but not least, while the stability analysis of random ecological systems offers powerful insights for the study of ecological dynamics and the understanding of how species interactions shape the stability and diversity of ecological communities, empirically applying the theory to real ecological systems poses a formidable challenge [79, 89, 126, 176, 177, 274, 275]. Notably, Jacquet et al. performed the local stability analysis of 116 real food webs sampled worldwide from marine, freshwater, and terrestrial habitats using the GLV model [126]. The community matrix M was constructed by multiplying the interaction matrix A with species biomass for each food web. The interaction coefficients of A can be translated from the parameters of the corresponding Ecopath model, a trophic model that provides accurate representation of feeding interactions within food webs [71]. The authors then measured the stability of food webs using the real part of the dominant eigenvalue of the community matrix M. Using randomization tests, they found that negative correlation between interaction strengths with high frequency of weak interactions is a strong stabilizing property in real food webs, see Fig. 6. Their findings reveal that empirical food webs exhibit some non-random characteristics that lead to the absence of a complexity–stability relationship [126].
@@ 4 Sign stability analysis
The notion of sign stability is of particular interest for ecology, economics, chemistry, and engineering [73, 91, 131, 158, 164, 206, 272, 273]. In an ecological system, the community matrix ${\bf M}$ (or the interaction matrix A) associated with certain models (e.g., the GLV model) might be only known qualitatively in the sense that the signs of the elements $\textbf{M}_{ij}$ can be determined with reasonable confidence, but the actual magnitudes may be very difficult to determine, see Fig. 7. Such matrices are often referred to as sign matrices. A sign matrix ${\bf M}$ is called sign stable (or sign semi-stable) if each of its eigenvalues has negative real part (or non-positive real part, respectively) for all numerical matrices of the same sign pattern [130, 131, 167, 206, 268].
@@ 4.1 Characterizations of sign stable matrices
A sign matrix ${\bf M}\in\mathbb{R}^{S\times S}$ is sign semi-stable if and only if the following conditions are all satisfied: (i) $\textbf{M}_{ii}\leq 0$ for $i=1,2,\dots,S$; (ii) $\textbf{M}_{ij}\textbf{M}_{ji}\leq 0$ for $i,j=1,2,\dots,S$ and $i\neq j$; (iii) there exists no elementary cycle of length $k\geq 3$ in the digraph generated by M [167, 206]. It is also known that if $\textbf{M}_{ii}<0$ for $i=1,2,\dots,S$, then (ii) and (iii) are necessary and sufficient for ${\bf M}$ to be sign stable. Quirk and Ruppert further claimed that (i)-(iii) as well as (iv) $\textbf{M}_{ii}<0$ for some $i$ and (v) $\text{det}(\textbf{M})\neq 0$ are both necessary and sufficient for ${\bf M}$ to be sign stable [206]. However, this was proved to be incorrect later by Jeffries et al. [130]. Later on, Yamada demonstrated that such an exceptional case is very rare and proved that the conditions (i)-(v) proposed by Quirk and Ruppert is actually necessary and sufficient for a system to be generically sign stable, i.e., sign stable for almost all parameter values except for some pathological cases with measure zero [268].
@@
Figure 7: Examples of signed ecological networks and their corresponding sign community matrices. (a) Self-regulated species 1 forms a commensal relationship with each species within the prey-predator pair of 2 and 3. (b) Linear trophic chain in which each successive species is preyed upon the subsequent one and species 3 is self-regulated. This figure was redrawn from [158].
@@
Quirk and Ruppert’s conditions (i)-(v) can be translated into ecological terms to provide a comprehensive characterization of sign stable patterns [158]. According to (i) and (iv), a sign stable ecological system should not include self-promoting species and must have at least one species with self-regulation. Condition (ii) necessitates the absence of both competitive and mutualistic relations. Condition (iii) states that there are no closed directed cycles with more than two edges in the network structure. Finally, condition (v) requires that the digraph generated by M must include a specific number of non-overlapping directed cycles that encompass all the nodes.
@@
Jeffries et al. introduced two additional conditions, distinct from (iv) and (v), called the coloring and matching conditions [130]. The two conditions, combined with (i)-(iii), are necessary and sufficient for sign stability. Let $\mathcal{R}_{M}=\{i\text{ }|\text{ }\textbf{M}_{ii}\neq 0\}$. An $\mathcal{R}_{M}$-coloring of an undirected graph is a partition of its nodes into two sets, black and white, such that each node in $\mathcal{R}_{M}$ is black (one of which can be empty), no black node has exactly one white neighbor, and each white node has at least one white neighbor. The coloring condition then states that in every $\mathcal{R}_{M}$-coloring of the undirected graph generated by M, all nodes are black. Additionally, a $(\mathcal{V}\sim\mathcal{R}_{M})$-complete matching in an undirected graph is a set $\mathcal{M}$ of disjoint edges such that an exact cover of the node set can be obtained using the pairs in $\mathcal{M}$ and certain singletons from $\mathcal{R}_{\mathcal{M}}$. The matching condition then asserts that the undirected graph generated by M admits a $(V\sim\mathcal{R}_{M})$-complete matching. However, the ecological systems characterized by Quirk and Ruppert [206] or Jeffries et al. [130] are not readily observable in nature.
@@ 4.2 Applications to ecological systems
The notion of sign stability has been applied to various ecological systems through different approaches. Dambacher et al. introduced two qualitative metrics – weighted feedback and weighted determinants based on the Hurwitz criterion [77], which can be recast into two conditions: (i) the coefficients of the characteristic equation of M must have the same sign; (ii) the corresponding Hurwitz determinants must all be positive. The two metrics offer a practical mean to identify the relative degree to which stable parameter space can be constrained based on system structure and complexity. Remarkably, Haraldsson et al. utilized the weighted feedback and weighted determinants to investigate the sign stability of social-ecological systems, which is an important tool to understand human-nature relations [118].
Moreover, Allesina and Pascual studied the sign stability of random and empirical food webs through a fairly intuitive approach [10]. Given a community matrix M (either randomly generated or empirical), first determine the stability based on its eigenvalues. If it is stable (in terms of $\text{Re}[\lambda(\textbf{M})]\leq 0$), generate 100 matrices that have the same sign pattern with random magnitude. Lastly, measure the percentage of the random generated matrices that are stable. The percentage can tell whether the stability is due to a particular combination of coefficients or the sign pattern of the network itself. Based on this approach, the authors demonstrated that predator-prey interactions can promote stability, highly robust to perturbations of interaction strength, in real ecological systems [10].
@@ 5 Diagonal stability analysis
The notion of diagonal stability was first introduced by Volterra in the 1930s [261]. It has been particularly useful for the stability analysis of ecological systems and other networked systems [29, 34, 45, 50, 134, 135, 186, 246]. A matrix A is called diagonally stable if there exists a diagonal matrix ${\bf P}\succ 0$ that renders
@@
We use the notation $\mathcal{D}(\textbf{P})$ for the class of diagonally stable matrices. The positive diagonal matrix P is often called Volterra multiplier in literature [210]. In many cases, the necessary and sufficient conditions for the Lyapunov stability of nonlinear systems are also the necessary and sufficient conditions for the diagonal stability of a certain matrix associated to the nonlinear system. This matrix naturally captures the underlying network structure of the nonlinear dynamical system.
Diagonal stability has been successfully applied to various types of dynamical systems [8, 140, 162, 243, 245, 267]. It has nice “structural consequences.” For example, the principal sub-matrices of a diagonally stable matrix are also diagonally stable, implying that all the corresponding “principal sub-systems” of a given diagonally stable system are diagonally stable [135]. In this section, we begin by introducing the theoretical framework for determining the Lyapunov stability of the GLV model via diagonal stability analysis. Subsequently, we delve into the characterizations of diagonal stability of the interaction matrix that is associated with special interconnection or network structures, which offer an effective mean of determining the stability of ecological systems.

### 07 Feb 2024 v2

@@ 3.2.2 Impacts of extinction
Extinction is generic in the GLV model with random interactions baron2023breakdown ; emary2021can ; pettersson2020stability . Recent studies have highlighted that an equilibrium point of the GLV model is locally stable if and only if all the eigenvalues of the reduced interaction matrix (i.e., the interaction matrix between the species in the surviving sub-community) have negative real parts barbier2021fingerprints ; baron2023breakdown ; biroli2018marginally ; stone2018feasibility . Baron et al. focused on the eigenvalue spectrum of the reduced interaction matrix in the GLV model, where the spectrum of the reduced interaction matrix also consists of a bulk set of eigenvalues and an outlier baron2023breakdown . Importantly, the authors demonstrated that the universality principle holds for the bulk region (i.e., following Ellipse Law) but not for the outlier eigenvalue. The outlier eigenvalue can be solved from the generating functional approach. Prediction of extinction boundary and new equilibrium after a primary extinction were also discussed in pettersson2020stability and emary2021can , respectively.
Figure 5: Eigenvalue distributions with time delay, where the blue teardrop-shaped area represents the stability region with time delay. (a) The eigenvalues (represented by red dots) are within the teardrop-shaped area, indicating that the corresponding GLV model with time delay is locally stable. (b) The eigenvalues are outside the teardrop-shaped area, indicating that the corresponding GLV model with time delay is unstable. $\tau=1$. The figure was redrawn from yang2023time .
@@ 3.2.3 Impacts of time delay
Similar to Pigani et al.’s work as discussed in Section 3.1.7, Yang et al. investigated the stability of the time-delayed GLV model yang2023time , which is defined as
@@
where $\textbf{M}_{\text{delay}}=\textbf{X}\textbf{A}$ is the community matrix with delay. For simplicity, X is set to I such that each species has unit abundance. However, the validation of this assumption remains an open problem yang2023time . They proved that (12) is stable if all the roots of the characteristic equation, defined as $H(z)=z-\lambda_{\text{delay}}e^{-z\tau}$ where $\lambda_{\text{delay}}$ are the eigenvalues of $\textbf{M}_{\text{delay}}$, have negative real parts. This equation implies that all the eigenvalues of $\textbf{M}_{\text{delay}}$ are required to be located in a teardrop-shaped region defined by $\text{Re}[z]<0$ to ensure stability, see Figure 5. Since the distribution of the eigenvalues of $\textbf{M}_{\text{delay}}$ (i.e., the interaction matrix A is this case) with random, mutualistic, or competitive interactions is well-established, it becomes straightforward to estimate $\max{(\text{Re}[z])}$. Based on the theoretical findings, the authors found that time delay plays a critical role in ecological systems, where large delay is often destabilizing, but short delay can considerably enhance community stability ($\tau=0.2$ and 0.5 were used for the small and large time delay in their simulations, respectively) yang2023time .
Similar conclusions were also obtained by Saeedian et al. when investigating the impact of time delay on the emergent stability patterns within the time-delayed GLV model (11) saeedian2022effect . Significantly, they further determined the existence of a Hopf bifurcation (i.e., the two complex conjugate eigenvalues of the community matrix, with non-zero imaginary part, simultaneously cross the imaginary axis into the right half-plane) at $\tau_{c}$, which is computed as
@@ 3.2.4 Impacts of stochastic noises
Real ecological systems are inherently stochastic with constant external perturbations and internal fluctuations tian2020estimating . Krumbeck et al. developed a theoretical framework for the analysis of temporal stability of ecological systems with stochastic noises krumbeck2021fluctuation . They provided analytical predictions for the power spectral density of the stochastic GLV model defined as
@@
where $i$ here is the imaginary number. They showed that different network structures (random, mutualistic, and competitive) have unique signatures in the spectrum of fluctuations. They further investigated the effect of trophic structures and identified a gap in the power spectral density, which indicates that high-level trophic structures contribute to enhanced long-term temporal stability krumbeck2021fluctuation .
@@ 3.2.5 Impacts of evolved system size
Previous work on the linear stability of the GLV model is concerned with ecological systems of predetermined and fixed sizes. Galla explored the stability of ecological systems with evolved system size using methods from statistical mechanics and the theory of disordered systems galla2018dynamically ; martin1973statistical ; mezard1987spin . Specifically, he exploited generating functionals to derive the effective dynamics for the GLV model (with $r_{i}=1$) computed as
@@
This result implies that predator-prey relationships enhance stability, while variability in species interactions promotes instability galla2018dynamically , which aligns with prior findings as reported in tang2014correlation ; yan2017degree .
Poley et al. applied Galla’s approach to analyze the stability of the GLV model with hierarchical interactions poley2023generalized defined as
@@
for $i=1,2,\dots,S^{a}$, where $B$ is the total number of local systems, and $a,b=1,2,\dots,B$ denote the indices of the local systems with the associated system size $S^{a}$ and $S^{b}$. Their findings indicate that a strong hierarchical structure is stabilizing but reduces diversity within the effective dynamics of the GLV model poley2023generalized . Similarly, Sidhom and Galla adapted Galla’s approach to study the GLV model with nonlinear feedback sidhom2020ecological defined as
@@
for $i=1,2,\dots,S$, where the function $g(u)=\frac{2au}{a+2|u|}$ denotes the nonlinear feedback ($a$ is the saturation parameter). This form of feedback was initially introduced to describe the predator’s growth rate when interacting with prey. It is plausible that the benefits from extra prey will eventually saturate as prey numbers become large. The authors concluded that the stability and diversity of ecological systems improves with the introduction of nonlinear feedback sidhom2020ecological .
Figure 6: Stability of real and permuted food webs in relation to complexity. These permuted food webs were constructed from real food webs by removing some non-random features, namely the row structure, topology, pairwise correlation, and interaction strength. The more the “rightmost”eigenvalues, i.e., $\text{Re}[\lambda_{1}]$, close to zero, the more stable of the food webs (because the diagonal elements of the community matrix are set to zero). This figure was redrawn from jacquet2016no .
@@ 3.2.6 Applications to real ecological systems
Last but not least, while the stability analysis of random ecological systems offers powerful insights for the study of ecological dynamics and the understanding of how species interactions shape the stability and diversity of ecological communities, empirically applying the theory to real ecological systems poses a formidable challenge de1995energetics ; emmerson2004predator ; jacquet2016no ; neutel2002stability ; neutel2007reconciling ; yodzis1981stability ; yonatan2022complexity . Notably, Jacquet et al. performed the local stability analysis of 116 real food webs sampled worldwide from marine, freshwater, and terrestrial habitats using the GLV model jacquet2016no . The community matrix M was constructed by multiplying the interaction matrix A with species biomass for each food web. The interaction coefficients of A can be translated from the parameters of the corresponding Ecopath model, a trophic model that provides accurate representation of feeding interactions within food webs christensen1992ecopath . The authors then measured the stability of food webs using the real part of the dominant eigenvalue of the community matrix M. Using randomization tests, they found that negative correlation between interaction strengths with high frequency of weak interactions is a strong stabilizing property in real food webs, see Figure 6. Their findings reveal that empirical food webs exhibit some non-random characteristics that lead to the absence of a complexity–stability relationship jacquet2016no .
@@ 4 Sign stability analysis
The notion of sign stability is of particular interest for ecology, economics, chemistry, and engineering clarke1975theorems ; eschenbach1998real ; jeffries1977matrix ; logofet1982sign ; may1973qualitative ; quirk1965qualitative ; yedavalli2008ecological ; yedavalli2009qualitative . In an ecological system, the community matrix ${\bf M}$ (or the interaction matrix A) associated with certain models (e.g., the GLV model) might be only known qualitatively in the sense that the signs of the elements $\textbf{M}_{ij}$ can be determined with reasonable confidence, but the actual magnitudes may be very difficult to determine, see Figure 7. Such matrices are often referred to as sign matrices. A sign matrix ${\bf M}$ is called sign stable (or sign semi-stable) if each of its eigenvalues has negative real part (or non-positive real part, respectively) for all numerical matrices of the same sign pattern Jeffries-74 ; jeffries1977matrix ; Maybee-69 ; quirk1965qualitative ; Yamada-Networks-90 .
@@ 4.1 Characterizations of sign stable matrices
A sign matrix ${\bf M}\in\mathbb{R}^{S\times S}$ is sign semi-stable if and only if the following conditions are all satisfied: (i) $\textbf{M}_{ii}\leq 0$ for $i=1,2,\dots,S$; (ii) $\textbf{M}_{ij}\textbf{M}_{ji}\leq 0$ for $i,j=1,2,\dots,S$ and $i\neq j$; (iii) there exists no elementary cycle of length $k\geq 3$ in the digraph generated by M Maybee-69 ; quirk1965qualitative . It is also known that if $\textbf{M}_{ii}<0$ for $i=1,2,\dots,S$, then (ii) and (iii) are necessary and sufficient for ${\bf M}$ to be sign stable. Quirk and Ruppert further claimed that (i)-(iii) as well as (iv) $\textbf{M}_{ii}<0$ for some $i$ and (v) $\text{det}(\textbf{M})\neq 0$ are both necessary and sufficient for ${\bf M}$ to be sign stable quirk1965qualitative . However, this was proved to be incorrect later by Jeffries et al. Jeffries-74 . Later on, Yamada demonstrated that such an exceptional case is very rare and proved that the conditions (i)-(v) proposed by Quirk and Ruppert is actually necessary and sufficient for a system to be generically sign stable, i.e., sign stable for almost all parameter values except for some pathological cases with measure zero Yamada-Networks-90 .
@@
Figure 7: Examples of signed ecological networks and their corresponding sign community matrices. (a) Self-regulated species 1 forms a commensal relationship with each species within the prey-predator pair of 2 and 3. (b) Linear trophic chain in which each successive species is preyed upon the subsequent one and species 3 is self-regulated. This figure was redrawn from logofet1982sign .
@@
Quirk and Ruppert’s conditions (i)-(v) can be translated into ecological terms to provide a comprehensive characterization of sign stable patterns logofet1982sign . According to (i) and (iv), a sign stable ecological system should not include self-promoting species and must have at least one species with self-regulation. Condition (ii) necessitates the absence of both competitive and mutualistic relations. Condition (iii) states that there are no closed directed cycles with more than two edges in the network structure. Finally, condition (v) requires that the digraph generated by M must include a specific number of non-overlapping directed cycles that encompass all the nodes.
@@
Jeffries et al. introduced two additional conditions, distinct from (iv) and (v), called the coloring and matching conditions Jeffries-74 . The two conditions, combined with (i)-(iii), are necessary and sufficient for sign stability. Let $\mathcal{R}_{M}=\{i\text{ }|\text{ }\textbf{M}_{ii}\neq 0\}$. An $\mathcal{R}_{M}$-coloring of an undirected graph is a partition of its nodes into two sets, black and white, such that each node in $\mathcal{R}_{M}$ is black (one of which can be empty), no black node has exactly one white neighbor, and each white node has at least one white neighbor. The coloring condition then states that in every $\mathcal{R}_{M}$-coloring of the undirected graph generated by M, all nodes are black. Additionally, a $(\mathcal{V}\sim\mathcal{R}_{M})$-complete matching in an undirected graph is a set $\mathcal{M}$ of disjoint edges such that an exact cover of the node set can be obtained using the pairs in $\mathcal{M}$ and certain singletons from $\mathcal{R}_{\mathcal{M}}$. The matching condition then asserts that the undirected graph generated by M admits a $(V\sim\mathcal{R}_{M})$-complete matching. However, the ecological systems characterized by Quirk and Ruppert quirk1965qualitative or Jeffries et al. Jeffries-74 are not readily observable in nature.
@@ 4.2 Applications to ecological systems
The notion of sign stability has been applied to various ecological systems through different approaches. Dambacher et al. introduced two qualitative metrics – weighted feedback and weighted determinants based on the Hurwitz criterion dambacher2003qualitative , which can be recast into two conditions: (i) the coefficients of the characteristic equation of M must have the same sign; (ii) the corresponding Hurwitz determinants must all be positive. The two metrics offer a practical mean to identify the relative degree to which stable parameter space can be constrained based on system structure and complexity. Remarkably, Haraldsson et al. utilized the weighted feedback and weighted determinants to investigate the sign stability of social-ecological systems, which is an important tool to understand human-nature relations haraldsson2020model .
Moreover, Allesina and Pascual studied the sign stability of random and empirical food webs through a fairly intuitive approach allesina2008network . Given a community matrix M (either randomly generated or empirical), first determine the stability based on its eigenvalues. If it is stable (in terms of $\text{Re}[\lambda(\textbf{M})]\leq 0$), generate 100 matrices that have the same sign pattern with random magnitude. Lastly, measure the percentage of the random generated matrices that are stable. The percentage can tell whether the stability is due to a particular combination of coefficients or the sign pattern of the network itself. Based on this approach, the authors demonstrated that predator-prey interactions can promote stability, highly robust to perturbations of interaction strength, in real ecological systems allesina2008network .
@@ 5 Diagonal stability analysis
The notion of diagonal stability was first introduced by Volterra in the 1930s volterra1959theory . It has been particularly useful for the stability analysis of ecological systems and other networked systems ballantine1970stabilization ; barker1978positive ; berman1983matrix ; bhaya1993discrete ; kaszkurewicz1996diagonal ; Kaszkurewicz-Book-00 ; pastravanu2010diagonal ; sun2023gallery . A matrix A is called diagonally stable if there exists a diagonal matrix ${\bf P}\succ 0$ that renders
@@
We use the notation $\mathcal{D}(\textbf{P})$ for the class of diagonally stable matrices. The positive diagonal matrix P is often called Volterra multiplier in literature Redheffer-SIAM-85 . In many cases, the necessary and sufficient conditions for the Lyapunov stability of nonlinear systems are also the necessary and sufficient conditions for the diagonal stability of a certain matrix associated to the nonlinear system. This matrix naturally captures the underlying network structure of the nonlinear dynamical system.
Diagonal stability has been successfully applied to various types of dynamical systems aleksandrov2018diagonal ; khong2016diagonal ; mason2006simultaneous ; sootla2017block ; sun2019diagonal ; wu2021diagonal . It has nice “structural consequences.” For example, the principal sub-matrices of a diagonally stable matrix are also diagonally stable, implying that all the corresponding “principal sub-systems” of a given diagonally stable system are diagonally stable Kaszkurewicz-Book-00 . In this section, we begin by introducing the theoretical framework for determining the Lyapunov stability of the GLV model via diagonal stability analysis. Subsequently, we delve into the characterizations of diagonal stability of the interaction matrix that is associated with special interconnection or network structures, which offer an effective mean of determining the stability of ecological systems.